OpenStudy (anonymous):

Line AC is drawn such that point B is between point A and point C. If AB = 2x, BC = 3x and AC = 370, then find the value of x and the lengths of AB and BC. x = AB = units BC = units

OpenStudy (anonymous):

2x+3x=370 5x=370 x=370/5=74 AB=2x=2(74)=148 BC=3x=3(74)=222

OpenStudy (anonymous):

PAKISTAN!!!

OpenStudy (anonymous):

uhh school

OpenStudy (anonymous):

im thinking of completing my studies in america hopefully

OpenStudy (anonymous):

i think so yes

OpenStudy (anonymous):

goodd..i like them both...math generallyy..

OpenStudy (anonymous):

i havent planned that far as of now ..but any suggestions would be great

OpenStudy (anonymous):

Well what kind of school do u want to go to?

OpenStudy (anonymous):

MATH..engineering...architecture

OpenStudy (anonymous):

Do u know what your GPA is?

OpenStudy (anonymous):

nah...

OpenStudy (anonymous):

but i got 100/100 in math the past exam

OpenStudy (anonymous):

i get a 90+ average...

OpenStudy (anonymous):

wud u recommend a nice place to live in america

OpenStudy (anonymous):

cooll...

OpenStudy (anonymous):

hot...i cant stand the cold

OpenStudy (anonymous):

hmmm...my brother is currently studying in waltham..how is that for a place to live in imean is it good

OpenStudy (anonymous):

and he told me how cold it is

OpenStudy (anonymous):

okwait lemme check

OpenStudy (anonymous):

@mohammad.arqum Did u get it?